# Joint identification of spatially variable genes via a network-assisted Bayesian regularization approach

### v1

@@ 2.4 Bayesian Inference
=$=\left\{\prod_{\left\{(i,j):r_{ij}=0\right\}}\operatorname{NB}\left(Y_{ij}\mid\mu_{ij},\phi\right)\right\}\times\mathrm{N}\left(\boldsymbol{\gamma}\mid\mathbf{0},\sigma_{\gamma}^{2}\left(\mathbf{L}+\varepsilon\mathbf{I}_{2p}\right)^{-1}\right)\times\operatorname{IG}\left(\sigma_{\boldsymbol{\gamma}}\mid a_{\gamma},b_{\gamma}\right)$
×$\quad\times\operatorname{Unif}\left(\lambda\mid\lambda_{l},\lambda_{u}\right)\times\mathrm{N}\left(\boldsymbol{\mu}_{0}\mid\mathbf{0},\operatorname{diag}\left(\sigma_{\mu_{01}}^{2},\ldots,\sigma_{\mu_{0p}}^{2}\right)\right)\times\mathrm{N}\left(\boldsymbol{\alpha}\mid\mathbf{0},\operatorname{diag}\left(\sigma_{\alpha_{1}}^{2},\ldots,\sigma_{\alpha_{K}}^{2}\right)\right)$
×$\quad\times\operatorname{Ga}\left(\phi\mid a_{\phi},b_{\phi}\right)\times\prod_{i=1}^{n}\prod_{j=1}^{p}\operatorname{Be-Bern}\left(r_{ij}\mid a_{\pi},b_{\pi}\right),$
The posterior sampling is conducted based on the MCMC algorithm. We first introduce the sampling variances $\tau_{\mu_{0}}^{2},\tau_{\alpha}^{2},\tau_{\phi}^{2},\tau_{\gamma}^{2},$ and $\tau_{\lambda}^{2}$ for $\mu_{0j}$’s, $\alpha_{k}$’s, $\phi$, $\boldsymbol{\gamma}$, and $\lambda$, respectively, and then conduct the following steps at each MCMC iteration, where the symbol “$\boldsymbol{-}$” in the condition position denotes “the rest parameters”.
@@
    Sample $\lambda^{*}$ from $N\left(\lambda,\tau_{\lambda}^{2}\right)$ truncated at interval $[\lambda_{l},\lambda_{u}]$, and accept $\lambda^{*}$ with probability $\min\left\{1,\frac{N_{+}\left(\lambda\mid\lambda^{*},\lambda_{l},\lambda_{u},\tau_{\lambda}^{2}\right)\prod_{\left\{\left(i,j\right):r_{ij}=0\right\}}\text{NB}\left(Y_{ij}\mid\lambda^{*},\boldsymbol{-}\right)}{N_{+}\left(\lambda^{*}\mid\lambda,\lambda_{l},\lambda_{u},\tau_{\lambda}^{2}\right)\prod_{\left\{\left(i,j\right):r_{ij}=0\right\}}\text{NB}\left(Y_{ij}\mid\lambda,\boldsymbol{-}\right)}\right\}.$
@@
Here, updating $r_{ij}$’s and $\sigma_{\boldsymbol{\gamma}}^{2}$ is achieved through the Gibbs sampler, while the Metropolis-Hasting (MH) algorithm is adopted for sampling $\mu_{0j}$’s, $\alpha_{k}$’s, $\phi$, and $\lambda$. For sampling $\boldsymbol{\gamma}$, we resort to the preconditioned Crank-Nicolson Langevin dynamics (pCNLD), which takes advantage of the gradient information of the target distribution to speed up convergence. Furthermore, with (5), pCNLD explicitly incorporates the network dependency into the sampling process. The details of the proposed pCNLD sampling are given in Supplementary Section S1. For $\rho_{j}^{(d)}$’s, since biological networks are often composed of multiple disconnected sub-networks, we propose adopting a set of group-wise weights for the $S$ sub-networks for more effectively utilizing network information. Specifically, a series of data-driven weights inversely proportional to the absolute effect sizes of the genes involved in the specific sub-networks are introduced, potentially facilitating the identification of SV genes with weak signals, which may be involved in the sub-networks consisting of SV genes with strong signals and small thresholds. The settings for the hyperparameters and sampling variances are provided in Supplementary Section S2.
@@
For the spatial modeling function $\mathcal{K}(\cdot)$, we follow the published studies and consider three most popular functions. Specifically, for $x_{id}~{}(d=1,2)$ which has been transformed to have mean 0 and standard deviation 1, we consider one linear function $\mathcal{K}_{1}(x_{id})=x_{id}$, two exponential functions $\mathcal{K}_{2}(x_{id})=\exp\left(-\frac{x_{id}}{2\left(l_{d}^{(1)}\right)^{2}}\right)$ and $\mathcal{K}_{3}(x_{id})=\exp\left(-\frac{x_{id}}{2\left(l_{d}^{(2)}\right)^{2}}\right)$, and two periodic functions $\mathcal{K}_{4}(x_{id})=\cos\left(\frac{2\pi x_{id}}{l_{d}^{(1)}}\right)$ and $\mathcal{K}_{5}(x_{id})=\cos\left(\frac{2\pi x_{id}}{l_{d}^{(2)}}\right)$, where $l_{d}^{(1)}$ and $l_{d}^{(2)}$ are the 40% and 60% quantiles of $|x_{1d}|,\cdots,|x_{nd}|$, respectively. To accommodate the fact that results may be sensitive to the choice of scale parameter, two values are considered for the exponential and periodic patterns. Then, we conduct the MCMC sampling for each $\mathcal{K}_{s}(\cdot)(s=1,\cdots,5)$ and obtain the estimated posterior expectation $\hat{\boldsymbol{\beta}}^{{\mathcal{K}_{s}}}=\frac{\sum_{m=1}^{M}\boldsymbol{\gamma}^{(m)}\boldsymbol{\mathbf{t}}_{\lambda^{(m)},\boldsymbol{\rho}^{(m)}}\left(\boldsymbol{\gamma}^{(m)}\right)}{\sum_{m=1}^{M}\boldsymbol{\mathbf{t}}_{\lambda^{(m)},\boldsymbol{\rho}^{(m)}}\left(\boldsymbol{\gamma}^{(m)}\right)},\hat{\boldsymbol{\mu}}_{0}^{{\mathcal{K}_{s}}}=\frac{\sum_{m=1}^{M}\boldsymbol{\mu}_{0}^{(m)}}{M},\hat{\boldsymbol{\alpha}}^{{\mathcal{K}_{s}}}=\frac{\sum_{m=1}^{M}\boldsymbol{\alpha}^{(m)}}{M}$, $\hat{\phi}^{{\mathcal{K}_{s}}}=\frac{\sum_{m=1}^{M}\phi^{(m)}}{M}$, and $\hat{\boldsymbol{R}}^{{\mathcal{K}_{s}}}=\frac{\sum_{m=1}^{M}\boldsymbol{R}^{(m)}}{M}$, where $\left\{\boldsymbol{\gamma}^{(m)},\lambda^{(m)},\boldsymbol{\rho}^{(m)},\boldsymbol{\mu}_{0}^{(m)},\boldsymbol{\alpha}^{(m)},\phi^{(m)},\boldsymbol{R}^{(m)}\right\}_{m=1}^{M}$ denotes the $M$ samples obtained after burn-in and thinning (we omit the dependence on $\mathcal{K}_{s}$ to simplify notation).
@@
To facilitate the combination of five models, instead of directly considering the values of $\hat{\boldsymbol{\beta}}^{{\mathcal{K}_{s}}}$ for SV gene identification, we further introduce a posterior inclusion probability vector estimated as ${\mathrm{PIP}}^{{\mathcal{K}_{s}}}=\frac{1}{M}\sum_{m=1}^{M}\boldsymbol{\mathbf{t}}_{\lambda^{(m)},\boldsymbol{\rho}^{(m)}}\left(\boldsymbol{\gamma}^{(m)}\right)$ and consider the posterior model probability $f\left(\mathcal{M}_{\mathcal{K}_{s}}\mid\mathbf{Y}\right)$ introduced in Quintana et al. (2011), where $\mathcal{M}_{\mathcal{K}_{s}}$ denotes the model with $\mathcal{K}_{s}(\cdot)$. Specifically, for each $\mathcal{M}_{\mathcal{K}_{s}}$, we calculate $f\left(\mathcal{M}_{\mathcal{K}_{s}}\mid\mathbf{Y}\right)=\frac{f\left(\mathbf{Y}\mid\mathcal{M}_{\mathcal{K}_{s}}\right)f\left(\mathcal{M}_{\mathcal{K}_{s}}\right)}{\sum_{s^{\prime}=1}^{5}f\left(\mathbf{Y}\mid\mathcal{M}_{\mathcal{K}_{s^{\prime}}}\right)f\left(\mathcal{M}_{\mathcal{K}_{s^{\prime}}}\right)}$. Here, $f\left(\mathcal{M}_{\mathcal{K}_{s}}\right)$ is the prior probability of model $\mathcal{M}_{\mathcal{K}_{s}}$, and we set a non-informative prior with $f\left(\mathcal{M}_{\mathcal{K}_{s}}\right)=1/5$. $f\left(\mathbf{Y}\mid\mathcal{M}_{\mathcal{K}_{s}}\right)$ is the likelihood function with the estimated parameters $\hat{\boldsymbol{\beta}}^{{\mathcal{K}_{s}}},\hat{\boldsymbol{\mu}}_{0}^{{\mathcal{K}_{s}}},\hat{\boldsymbol{\alpha}}^{{\mathcal{K}_{s}}},\hat{\phi}^{\mathcal{K}_{s}}$, and $\hat{\boldsymbol{R}}^{{\mathcal{K}_{s}}}$ under model $\mathcal{M}_{\mathcal{K}_{s}}$. We then utilize $f\left(\mathcal{M}_{\mathcal{K}_{s}}\mid\mathbf{Y}\right)$’s as the model-specific weights to obtain the combined $\operatorname{PIP}=\sum\limits_{s=1}^{5}f\left(\mathcal{M}_{\mathcal{K}_{s}}\mid\mathbf{Y}\right)\mathrm{PIP}^{{\mathcal{K}_{s}}}$. Moreover, as gene $j$ with at least one non-zero $\beta^{(d)}_{j}\left(d=1,2\right)$ is identified as SV, we introduce $\widetilde{\operatorname{PIP}}_{j}=\max\left(\operatorname{PIP}_{j},\mathrm{PIP}_{j+p}\right)$ for $j=1,\ldots,p$. Based on $\widetilde{\operatorname{PIP}}_{j}$’s, the Bayesian false discovery rate (BFDR) control strategy is adopted for controlling multiplicity. Specifically, $\operatorname{BFDR}(c)=\frac{\sum_{j=1}^{p}\left(1-\widetilde{\operatorname{PIP}}_{j}\right)\mathrm{I}\left(1-\widetilde{\operatorname{PIP}}_{j}<c\right)}{\sum_{j=1}^{p}\mathrm{I}\left(1-\widetilde{\operatorname{PIP}}_{j}<c\right)}$ with $\operatorname{BFDR}(c)$ being the desired significance level. We set $\operatorname{BFDR}(c)$ as 0.05 in our numerical studies. Then, the SV gene set is defined as $\left\{j:\widetilde{\operatorname{PIP}}_{j}\geq c\right\}$. To achieve improved stability and better false negative control, in numerical studies, we run five MCMC chains independently, and the genes identified in more than 80% of the chains are finally identified as SV.
@@
To improve computational efficiency, parallelization is implemented with the R package RcppParallel. Specifically, the marginal sequential sampling is divided into parallel programming to reduce computer time. For $\boldsymbol{\gamma}$ with the dependency Laplacian matrix incorporated, benefiting from the sparse and block-wise nature of biological networks, the parallelism block-wise sampler is adopted to avoid sampling from a high-dimensional multivariate normal distribution, which further accelerates computation. Take a simulated dataset with $p=5,000$ and $n=1,024$ as an example. The average computer time is about 1.01 minutes for 100 iterations on a MacBook Pro with an Intel Core i5 CPU and 16GB RAM.
@@ 3.1 Basic Simulations
Simulation studies are conducted under the following settings. (a) $n=1,024$ spots located on a 32 by 32 square lattice, $p=5,000$ genes, and $K=6$ cell types. (b) The square lattice is partitioned into three regions as displayed in Figure S1, where the cellular compositions $\boldsymbol{w}_{i}$’s are independently sampled from Dirichlet distributions $\operatorname{Dirc}(1,1,1,1,1,1)$ (Region 1), $\operatorname{Dirc}(3,5,7,9,11,13)$ (Region 2), and $\operatorname{Dirc}(18,16,14,12,10,8)$ (Region 3). (c) Consider three types of spatial pattern $\mathcal{K}\left(x_{id}\right)$, including $x_{id}$ (Linear), $\exp\left(-\frac{x_{id}}{2}\right)$ (Exponential), and $\cos\left(2\pi x_{id}\right)$ (Periodic). (d) The networks are block-wise and composed of 100 disconnected sub-networks with 50 nodes each. For each sub-network, two types of network structure, namely Star and Scale-free, are considered to mimic the real-world transcription factor regulatory network and interaction network with scale-free properties, respectively. Illustrative examples of the sub-networks are presented in Figure S2(A) and S2(B). (e) All genes in the first ten sub-networks are SV, leading to a total of 500 SV genes. Both positive and negative signals are considered with various levels of magnitude. More detailed settings are provided in Supplementary Section S3.1. (f) The spatial transcriptomics count data is generated from model (1) with the dispersion parameter $\phi$ being 10. Two settings of dropout rates, 0.1 and 0.5, are considered, representing low and high sparsity. The baseline parameter ${\mu}_{0j}$ and cell-type-specific effect ${\alpha}_{k}$ are independently generated from $N(2,0.5^{2})$ and $N(0,3.5^{2})$, respectively. There are 12 scenarios, comprehensively covering a wide spectrum with different patterns of spatial expressions, different structures of networks and the corresponding spatially variable signals, and different degrees of sparsity.
@@
In addition to the proposed approach, six alternatives are considered: (a) SpatialDE (Svensson et al., 2018), a likelihood ratio test method based on Gaussian process regression, (b) SPARK (Sun et al., 2020), a method based on a Poisson log linear model with a Gaussian process incorporated, (c) SPARK-X (Zhu et al., 2021), a scalable non-parametric test built on a robust covariance test framework, (d) MERINGUE (Miller et al., 2021), a test based on spatial cross-correlation analysis, (e) CTSV and CTSV-g (Yu and Luo, 2022), a test based on the ZINB model for cell-type-specific SV gene identification and global SV gene identification, respectively, and (f) HEARTSVG (Yuan et al., 2023), a distribution-free method that employs the exclusion of non SV genes to infer the presence of SV genes. Among them, SpatialDE, SPARK, CTSV, and CTSV-g are parametric, while SPARK-X, MERINGUE, and HEARTSVG are non-parametric. All the alternatives conduct tests marginally and implement multiple testing control. The implementation details for the competing methods are given in Supplementary Section S3.2.
@@
Table 1: Simulation results under the scenarios with a low dropout rate, where FDR (BFDR) is controlled to be $<$0.05. In each cell, mean (SD) is based on 50 replicates.

### v2

@@ 2.4 Bayesian Inference
$\int f\left(\mathbf{Y}\mid\boldsymbol{\gamma},\boldsymbol{R},\boldsymbol{\mu}_{0},\boldsymbol{\alpha},\phi,\lambda\right)f\left(\boldsymbol{\gamma}\mid\sigma_{\gamma}^{2}\right)f\left(\sigma_{\gamma}^{2}\right)f(\lambda)f\left(\boldsymbol{\mu}_{0}\right)f\left(\boldsymbol{\alpha}\right)f(\phi)f(\boldsymbol{R}\mid\boldsymbol{\pi})f(\boldsymbol{\pi})d\boldsymbol{\pi}$
$=\left\{\prod_{\left\{(i,j):r_{ij}=0\right\}}\operatorname{NB}\left(Y_{ij}\mid\mu_{ij},\phi\right)\right\}\times\mathrm{N}\left(\boldsymbol{\gamma}\mid\mathbf{0},\sigma_{\gamma}^{2}\left(\mathbf{L}+\varepsilon\mathbf{I}_{2p}\right)^{-1}\right)\times\operatorname{IG}\left(\sigma_{\boldsymbol{\gamma}}\mid a_{\gamma},b_{\gamma}\right)$
$\quad\times\operatorname{Unif}\left(\lambda\mid\lambda_{l},\lambda_{u}\right)\times\mathrm{N}\left(\boldsymbol{\mu}_{0}\mid\mathbf{0},\operatorname{diag}\left(\sigma_{\mu_{01}}^{2},\ldots,\sigma_{\mu_{0p}}^{2}\right)\right)\times\mathrm{N}\left(\boldsymbol{\alpha}\mid\mathbf{0},\operatorname{diag}\left(\sigma_{\alpha_{1}}^{2},\ldots,\sigma_{\alpha_{K}}^{2}\right)\right)$
$\quad\times\operatorname{Ga}\left(\phi\mid a_{\phi},b_{\phi}\right)\times\prod_{i=1}^{n}\prod_{j=1}^{p}\operatorname{Be-Bern}\left(r_{ij}\mid a_{\pi},b_{\pi}\right),$
@@
    Sample $\lambda^{*}$ from $N\left(\lambda,\tau_{\lambda}^{2}\right)$ truncated at interval $\lambda_{l},\lambda_{u}]$, and accept $\lambda^{*}$ with probability $\min\left\{1,\frac{N_{+}\left(\lambda\mid\lambda^{*},\lambda_{l},\lambda_{u},\tau_{\lambda}^{2}\right)\prod_{\left\{\left(i,j\right):r_{ij}=0\right\}}\text{NB}\left(Y_{ij}\mid\lambda^{*},\boldsymbol{-}\right)}{N_{+}\left(\lambda^{*}\mid\lambda,\lambda_{l},\lambda_{u},\tau_{\lambda}^{2}\right)\prod_{\left\{\left(i,j\right):r_{ij}=0\right\}}\text{NB}\left(Y_{ij}\mid\lambda,\boldsymbol{-}\right)}\right\}.$
@@
Here, updating $r_{ij}$’s and $\sigma_{\boldsymbol{\gamma}}^{2}$ is achieved through the Gibbs sampler, while the Metropolis-Hasting (MH) algorithm is adopted for sampling $\mu_{0j}$’s, $\alpha_{k}$’s, $\phi$, and $\lambda$. For sampling $\boldsymbol{\gamma}$, we resort to the preconditioned Crank-Nicolson Langevin dynamics (pCNLD), which takes advantage of the gradient information of the target distribution to speed up convergence. Furthermore, with (5), pCNLD explicitly incorporates the network dependency into the sampling process. The details of the proposed pCNLD sampling are given in Supplementary Section S1 (Wu et al., 2025). For $\rho_{j}^{(d)}$’s, since biological networks are often composed of multiple disconnected sub-networks, we propose adopting a set of group-wise weights for the $S$ sub-networks for more effectively utilizing network information. Specifically, a series of data-driven weights inversely proportional to the absolute effect sizes of the genes involved in the specific sub-networks are introduced, potentially facilitating the identification of SV genes with weak signals, which may be involved in the sub-networks consisting of SV genes with strong signals and small thresholds. The settings for the hyperparameters and sampling variances are provided in Supplementary Section S2 (Wu et al., 2025).
@@
For the spatial modeling function $\mathcal{K}(x_{id})$, as discussed above, we consider three most popular functions as recommended in the published studies. Specifically, for $x_{id}~(d=1,2)$ which has been transformed to have mean 0 and standard deviation 1, we consider one linear function $\mathcal{K}_{1}(x_{id})=x_{id}$, two exponential functions $\mathcal{K}_{2}(x_{id})=\exp\left(-\frac{x_{id}}{2\left(l_{d}^{(1)}\right)^{2}}\right)$ and $\mathcal{K}_{3}(x_{id})=\exp\left(-\frac{x_{id}}{2\left(l_{d}^{(2)}\right)^{2}}\right)$, and two periodic functions $\mathcal{K}_{4}(x_{id})=\cos\left(\frac{2\pi x_{id}}{l_{d}^{(1)}}\right)$ and $\mathcal{K}_{5}(x_{id})=\cos\left(\frac{2\pi x_{id}}{l_{d}^{(2)}}\right)$, where $l_{d}^{(1)}$ and $l_{d}^{(2)}$ are the 40% and 60% quantiles of $|x_{1d}|,\cdots,|x_{nd}|$, respectively. To accommodate the fact that results may be sensitive to the choice of scale parameter, two values are considered for the exponential and periodic patterns. Then, we conduct the MCMC sampling for each $\mathcal{K}_{s}(\cdot)(s=1,\cdots,5)$ and obtain the estimated posterior expectation $\hat{\boldsymbol{\beta}}^{{\mathcal{K}_{s}}}=\frac{\sum_{m=1}^{M}\boldsymbol{\gamma}^{(m)}\boldsymbol{\mathbf{t}}_{\lambda^{(m)},\boldsymbol{\rho}^{(m)}}\left(\boldsymbol{\gamma}^{(m)}\right)}{\sum_{m=1}^{M}\boldsymbol{\mathbf{t}}_{\lambda^{(m)},\boldsymbol{\rho}^{(m)}}\left(\boldsymbol{\gamma}^{(m)}\right)},\hat{\boldsymbol{\mu}}_{0}^{{\mathcal{K}_{s}}}=\frac{\sum_{m=1}^{M}\boldsymbol{\mu}_{0}^{(m)}}{M},\hat{\boldsymbol{\alpha}}^{{\mathcal{K}_{s}}}=\frac{\sum_{m=1}^{M}\boldsymbol{\alpha}^{(m)}}{M}$, $\hat{\phi}^{{\mathcal{K}_{s}}}=\frac{\sum_{m=1}^{M}\phi^{(m)}}{M}$, and $\hat{\boldsymbol{R}}^{{\mathcal{K}_{s}}}=\frac{\sum_{m=1}^{M}\boldsymbol{R}^{(m)}}{M}$, where $\left\{\boldsymbol{\gamma}^{(m)},\lambda^{(m)},\boldsymbol{\rho}^{(m)},\boldsymbol{\mu}_{0}^{(m)},\boldsymbol{\alpha}^{(m)},\phi^{(m)},\boldsymbol{R}^{(m)}\right\}_{m=1}^{M}$ denotes the $M$ samples obtained after burn-in and thinning (we omit the dependence on $\mathcal{K}_{s}$ to simplify notation).
@@
To facilitate the combination of five models, instead of directly considering the values of $\hat{\boldsymbol{\beta}}^{{\mathcal{K}_{s}}}$ for SV gene identification, we further introduce a posterior inclusion probability vector estimated as ${\mathrm{PIP}}^{{\mathcal{K}_{s}}}=\frac{1}{M}\sum_{m=1}^{M}\boldsymbol{\mathbf{t}}_{\lambda^{(m)},\boldsymbol{\rho}^{(m)}}\left(\boldsymbol{\gamma}^{(m)}\right)$ and consider the posterior model probability $f\left(\mathcal{M}_{\mathcal{K}_{s}}\mid\mathbf{Y}\right)$ introduced in Quintana et al. (2011), where $\mathcal{M}_{\mathcal{K}_{s}}$ denotes the model with $\mathcal{K}_{s}(\cdot)$. Specifically, for each $\mathcal{M}_{\mathcal{K}_{s}}$, we calculate $f\left(\mathcal{M}_{\mathcal{K}_{s}}\mid\mathbf{Y}\right)=\frac{f\left(\mathbf{Y}\mid\mathcal{M}_{\mathcal{K}_{s}}\right)f\left(\mathcal{M}_{\mathcal{K}_{s}}\right)}{\sum_{s^{\prime}=1}^{5}f\left(\mathbf{Y}\mid\mathcal{M}_{\mathcal{K}_{s^{\prime}}}\right)f\left(\mathcal{M}_{\mathcal{K}_{s^{\prime}}}\right)}$. Here, $f\left(\mathcal{M}_{\mathcal{K}_{s}}\right)$ is the prior probability of model $\mathcal{M}_{\mathcal{K}_{s}}$, and we set a non-informative prior with $f\left(\mathcal{M}_{\mathcal{K}_{s}}\right)=1/5$. $f\left(\mathbf{Y}\mid\mathcal{M}_{\mathcal{K}_{s}}\right)$ is the likelihood function with the estimated parameters $\hat{\boldsymbol{\beta}}^{{\mathcal{K}_{s}}},\hat{\boldsymbol{\mu}}_{0}^{{\mathcal{K}_{s}}},\hat{\boldsymbol{\alpha}}^{{\mathcal{K}_{s}}},\hat{\phi}^{\mathcal{K}_{s}}$, and $\hat{\boldsymbol{R}}^{{\mathcal{K}_{s}}}$ under model $\mathcal{M}_{\mathcal{K}_{s}}$. We then utilize $f\left(\mathcal{M}_{\mathcal{K}_{s}}\mid\mathbf{Y}\right)$’s as the model-specific weights to obtain the combined $\operatorname{PIP}=\sum\limits_{s=1}^{5}f\left(\mathcal{M}_{\mathcal{K}_{s}}\mid\mathbf{Y}\right)\mathrm{PIP}^{{\mathcal{K}_{s}}}$. Such an ensemble strategy allows us to comprehensively consider the effects across different kernel functions to improve identification accuracy. Moreover, as gene $j$ with at least one non-zero $\beta^{(d)}_{j}\left(d=1,2\right)$ is identified as SV, we introduce $\widetilde{\operatorname{PIP}}_{j}=\max\left(\operatorname{PIP}_{j},\mathrm{PIP}_{j+p}\right)$ for $j=1,\ldots,p$. Based on $\widetilde{\operatorname{PIP}}_{j}$’s, the Bayesian false discovery rate (BFDR) control strategy is adopted for controlling multiplicity. Specifically, $\operatorname{BFDR}(c)=\frac{\sum_{j=1}^{p}\left(1-\widetilde{\operatorname{PIP}}_{j}\right)\mathrm{I}\left(1-\widetilde{\operatorname{PIP}}_{j}<c\right)}{\sum_{j=1}^{p}\mathrm{I}\left(1-\widetilde{\operatorname{PIP}}_{j}<c\right)}$ with $\operatorname{BFDR}(c)$ being the desired significance level. We set $\operatorname{BFDR}(c)$ as 0.05 in our numerical studies. Then, the SV gene set is defined as $\left\{j:\widetilde{\operatorname{PIP}}_{j}\geq c\right\}$. To achieve improved stability and better false negative control, in numerical studies, we run five MCMC chains independently, and the genes identified in more than 80% of the chains are finally identified as SV.
@@
To improve computational efficiency, parallelization is implemented with the R package RcppParallel. Specifically, the marginal sequential sampling is divided into parallel programming to reduce computer time. For $\boldsymbol{\gamma}$ with the dependency Laplacian matrix incorporated, benefiting from the sparse and block-wise nature of biological networks, the parallelism block-wise sampler is adopted to avoid sampling from a high-dimensional multivariate normal distribution, which further accelerates computation. More discussions on the computer time of the proposed algorithm are provided in Supplementary Section S3 (Wu et al., 2025).
@@ 3.1 Basic Simulations
Simulation studies are conducted under the following settings. (a) $n=1,024$ spots located on a 32 by 32 square lattice, $p=5,000$ genes, and $K=6$ cell types. (b) The square lattice is partitioned into three regions as displayed in Figure S1, where the cellular compositions $\boldsymbol{w}_{i}$’s are independently sampled from Dirichlet distributions $\operatorname{Dirc}(1,1,1,1,1,1)$ (Region 1), $\operatorname{Dirc}(3,5,7,9,11,13)$ (Region 2), and $\operatorname{Dirc}(18,16,14,12,10,8)$ (Region 3). (c) Consider three types of spatial pattern $\mathcal{K}\left(x_{id}\right)$, including $x_{id}$ (Linear), $\exp\left(-\frac{x_{id}}{2}\right)$ (Exponential), and $\cos\left(2\pi x_{id}\right)$ (Periodic). (d) The networks are block-wise and composed of 100 disconnected sub-networks with 50 nodes each. For each sub-network, two types of network structure, namely Star and Scale-free, are considered to mimic the real-world transcription factor regulatory network and interaction network with scale-free properties, respectively. Illustrative examples of the sub-networks are presented in Figure S2(A) and S2(B). (e) All genes in the first ten sub-networks are SV, leading to a total of 500 SV genes. Both positive and negative signals are considered with various levels of magnitude. More detailed settings are provided in Supplementary Section S4 (Wu et al., 2025). (f) The spatial transcriptomics count data is generated from model (1) with the dispersion parameter $\phi$ being 10. Two dropout rates settings, 0.1 and 0.5, are considered, representing low and high sparsity. The baseline parameter ${\mu}_{0j}$ and the cell-type-specific effect ${\alpha}_{k}$ are independently generated from $N(2,0.5^{2})$ and $N(0,3.5^{2})$, respectively. There are 12 scenarios (Table S1), comprehensively covering a wide spectrum with different patterns of spatial expressions, different structures of networks and the corresponding spatially variable signals, and different degrees of sparsity.
@@
In addition to the proposed approach, nine alternatives are considered. SpatialDE (Svensson, Teichmann and Stegle, 2018) is a likelihood ratio test method based on Gaussian process regression. SPARK (Sun, Zhu and Zhou, 2020) is a method built on a Poisson log-linear model with a Gaussian process incorporated. SPARK-X (Zhu, Sun and Zhou, 2021) is a scalable non-parametric test constructed on a robust covariance test framework. HRG (Wu et al., 2022) serves as a method to detect informative genes that exhibit regional expression patterns within the cell-cell similarity network. MERINGUE (Miller et al., 2021) is a graph-based testing method relying on spatial cross-correlation analysis. CTSV and CTSV-g (Yu and Luo, 2022) are tests based on the ZINB model, with CTSV used for cell-type-specific SV gene identification and CTSV-g for global SV gene identification, respectively. SINFONIA (Jiang et al., 2023) is a graph-based method that ranks genes according to spatial autocorrelation measurements. nnSVG (Weber et al., 2023) is a test based on nearest-neighbor Gaussian processes. HEARTSVG (Yuan et al., 2024) is a distribution-free method which employs the exclusion of non-SV genes to infer the presence of SV genes. Among these alternatives, SpatialDE, SPARK, CTSV, CTSV-g, and nnSVG are parametric, while SPARK-X, HRG, MERINGUE, SINFONIA, and HEARTSVG are non-parametric. In addition, HRG, MERINGUE, SINFONIA, and nnSVG adopt graph strategies. All the alternatives except HRG and SINFONIA conduct tests marginally and implement multiple testing control, whereas HRG and SINFONIA provide gene ranking. The implementation details for the competing methods are given in Supplementary Section S5 (Wu et al., 2025).
@@
Table 1: Simulation results under the scenarios with a low dropout rate, where FDR (BFDR) is controlled to be <<0.05. In each cell, mean (SD) is based on 50 replicates.